OpenStudy (jadebirdfly):

A housefly sits on the outer edge of a rotating circular ceiling fan with a diameter of 6 feet. The fan rotates constantly at a rate of 20 revolutions per minute. How many minutes had the housefly been on the fan during the time it took to travel 19,404π feet? Express your answer to the nearest whole number.

OpenStudy (dumbcow):

\[t = \frac{d}{r} = \frac{19,404 \pi}{20(6 \pi)} = 161.7\]